OpenStudy (anonymous):

steve subscribed to 10 podcasts for 340 minutes. he used his 2 favorite tags, Hobbies and Recreation and Soliloquies. Each of the hobbies and recreation lasted about 32 minutes. each soliloquies episode lasted 42 minutes. to how many of each tag did steve subscribe?

OpenStudy (texaschic101):

hobbies and recreation = h soliloquies = s h + s = 10 ---> h = 10 - s 32h + 42s = 340 now substitute 10 - s in for h in the second equation. In doing that, it should give you s. Then substitute what you got for s into the first equation and that should give you y. You can then check you answer by substituting what you got for s and h into either of the equations, and they should equal.If they don't equal, then there was a mistake made. Try it and see what you get :)

OpenStudy (anonymous):

okay so i still dont understand could you work out the steps maybe show me the work for the answers

OpenStudy (texaschic101):

yes I can.... do you understand how I got the equations ?

OpenStudy (anonymous):

yes i do i just need the work to understand the answer i need

OpenStudy (texaschic101):

h + s =10 ---> h = 10 - s 32h + 42s = 340 now sub 10 - s in for h in the second equation... 32 (10 - s) + 42s = 340 (now distribute 32 through the parenthesis) 320 - 32s + 42s = 340 (combine like terms) 320 + 10s = 340 (subtract 320 from both sides to get 10s by itself) 320 + 10s - 320 = 340 - 320 (combine like terms) 10s = 20 (divide both sides by 10 to get s by itself) 10s/10 = 20/10 s = 2 now sub 2 in for s in the first equation... h = 10 - s h = 10 - 2 h = 8 now sub them both in either of the equations to check it... h = 10 - s 8 = 10 - 2 8 = 8 (correct) or/and check it in the other equation... 32h + 42s = 340 32(8) + 42(2) = 340 256 + 84 = 340 340 = 340 (correct) answer : hobbies and recreation (h) = 8 soliloquies (s) = 2

OpenStudy (anonymous):

thank you sooooo much you are awesomeeee

OpenStudy (texaschic101):

your welcome....thats what I am here for :)